# Interior Spectral Windows and Transport for Discrete Fractional Laplacians on $d$-Dimensional Hypercubic Lattices

### v1

@@ 1. Introduction and main results
Classical discrete Laplacians capture only short-range, nearest-neighbour interactions. Modern models of anomalous diffusion and long-range transport, however, require operators whose influence decays algebraically rather than being compactly supported. This motivates the study of fractional powers of the discrete Laplacian s [JoKoPaSe, BaDeJiLi, Kw, GHPKLBMH] and their magnetic perturbations [FaZh]: they naturally describe long-range effects on graphs [TaZa], generate nonlocal transport phenomena [ScSo], and provide a bridge between discrete models and continuous nonlocal PDEs.
@@
The second approach is a Series or kernel representations. Explicit expansions in terms of lattice shifts or finite differences reveal how distant sites influence each other [Kw, Ba2022], a key point for both numerics and decay analysis. In the translation-invariant case of the infinite lattice $\mathbb{Z}^{d}$, the situation is considerably simpler. The discrete Laplacian is diagonalized by the Fourier transform, with symbol
@@
In what follows we focus on the fractional discrete anisotropic Laplacian $\Delta^{\vec{\mathbf{r}}}_{{\mathbb{Z}}^{d}}$ acting on $\ell^{2}({\mathbb{Z}}^{d})$, where $\vec{\mathbf{r}}=(r_{1},\dots,r_{d})\neq 0$:
$\Delta^{\vec{\mathbf{r}}}_{{\mathbb{Z}}^{d}}:=\sum_{j=1}^{d}\mathrm{id}^{\otimes(j-1)}\otimes\Delta_{{\mathbb{Z}}}^{\,r_{j}}\otimes\mathrm{id}^{\otimes(d-j)}.$
@@
$H:=\Delta^{\vec{\mathbf{r}}}_{{\mathbb{Z}}^{d}}+W(Q),$
@@
$\sigma\!\left(\Delta_{{\mathbb{Z}}^{d}}^{\vec{\mathbf{r}}}\right)=\Bigl{\{}\ \sum_{j=1}^{d}\lambda_{j}\ \Big{|}\ \lambda_{j}\in\sigma\!\left(\Delta_{{\mathbb{Z}}}^{\,r_{j}}\right)\ \Bigr{\}},$
@@
$\operatorname{Thr}_{\mathrm{fin}}\!\bigl{(}H_{{\mathbb{Z}}^{d}}^{\vec{\mathbf{r}}}\bigr{)}=\Bigl{\{}\ \sum_{j\in N}4^{\,r_{j}}+\sum_{j\in P}\epsilon_{j}\,4^{\,r_{j}}\ :\ \epsilon_{j}\in\{0,1\}\ \Bigr{\}},$
@@
Let $H=\Delta_{{\mathbb{Z}}^{d}}^{\vec{\mathrm{r}}}+W(Q)$ with real $W$ satisfying (H0) and (H1). Then $H\in\mathcal{C}^{1,1}_{\mathrm{loc}}(A_{{\mathbb{Z}}^{d}})$ on each $\mathcal{I}\Subset\sigma(H_{0})^{\circ}\setminus\operatorname{Thr}(\Delta_{{\mathbb{Z}}^{d}}^{\vec{\mathbf{r}}})$, and localized Mourre theory yields:
@@
This work develops a localized (nonlocal) Mourre theory for fractional discrete Laplacians $\Delta^{\vec{\mathbf{r}}}_{{\mathbb{Z}}^{d}}$ with anisotropic orders $\vec{\mathbf{r}}\neq 0$. Within this unified framework we prove, on compact energy windows away from thresholds, the Limiting Absorption Principle and asymptotic completeness for fractional difference operators. To the best of our knowledge, this is the first rigorous Mourre/LAP/scattering treatment for fractional discrete Laplacians.
@@ 2.1. Fractional powers of discrete Laplacians in the unidimensional setting
The shift operator acts on $\ell^{2}(\mathbb{Z})$, we will be denoted simply by $U$, and its adjoint is the backward shift $U^{*}$ given by $(U^{*}f)(n)=f(n+1)$. These satisfy $UU^{*}=U^{*}U=\mathrm{id}_{\ell^{2}(\mathbb{Z})}$, i.e., $U$ is unitary on $\ell^{2}(\mathbb{Z})$.
We also denote by $Q$ the position operator on $\ell^{2}(\mathbb{Z})$ defined by $(Qf)(n)=nf(n)$. For any $s\in\mathbb{R}$, we define the weighted space $\ell^{2}_{s}(\mathbb{Z}):=\{f\in\ell^{2}(\mathbb{Z})\mid\|\langle Q\rangle^{s}f\|<\infty\}$, where $\langle Q\rangle:=(1+Q^{2})^{1/2}$. Let $\Delta_{\mathbb{Z}}:=2\mathrm{id}-U-U^{*}$ denote the standard discrete Laplacian on $\ell^{2}(\mathbb{Z})$. This operator is self-adjoint and positive semi-definite on $\ell^{2}(\mathbb{Z})$. To extend the notion of fractional powers $\Delta^{r}_{\mathbb{Z}}$ for any real $r$, we use the discrete Fourier transform framework, interpreting the powers via spectral calculus in the space of tempered distributions. This allows us to define $\Delta^{r}_{\mathbb{Z}}$ even when $r$ is negative.
@@
$\widehat{\Delta_{{\mathbb{Z}}}^{\,r}f}(\theta)=\bigl{(}2(1-\cos\theta)\bigr{)}^{r}\,\widehat{f}(\theta),\qquad(\Delta_{{\mathbb{Z}}}^{\,r}u)(n)=\sum_{k\in{\mathbb{Z}}}a_{r}(k)\,u(n-k),$
@@
$a_{r}(k)=\frac{1}{2\pi}\int_{-\pi}^{\pi}\bigl{(}2(1-\cos\theta)\bigr{)}^{r}e^{{\rm i}k\theta}\,d\theta.$
@@
$(\mathcal{F}u)(\theta)=\frac{1}{\sqrt{2\pi}}\sum_{n\in{\mathbb{Z}}}u(n)e^{-{\rm i}n\theta}.$
@@
$\mathcal{F}\,U\,\mathcal{F}^{*}f(\theta:)=M_{e^{-{\rm i}\theta}}f(\theta)=:e^{-{\rm i}\theta}f(\theta),\qquad\mathcal{F}\,U^{*}\,\mathcal{F}^{*}=M_{e^{{\rm i}\theta}},$
@@
By the (Borel) functional calculus for the nonnegative self-adjoint operator $\Delta_{{\mathbb{Z}}}$,
$\mathcal{F}\,\Delta_{{\mathbb{Z}}}^{\,r}\,\mathcal{F}^{*}=M_{\vartheta^{r}},$
so for $u\in\ell^{2}({\mathbb{Z}})$ we have $\widehat{\Delta_{{\mathbb{Z}}}^{\,r}u}(\theta)=\vartheta(\theta)^{r}\,\widehat{u}(\theta)$. The Fourier transform extends to a topological isomorphism $\mathcal{F}:\mathcal{S}^{\prime}({\mathbb{Z}})\to\mathcal{S}^{\prime}(S^{1})$; by density/duality the same identity holds for all $f\in\mathcal{S}^{\prime}({\mathbb{Z}})$, proving the multiplier formula.
Let $a_{r}\in\mathcal{S}^{\prime}({\mathbb{Z}})$ be defined by its Fourier transform
$\widehat{a_{r}}(\theta)=\vartheta(\theta)^{r},\qquad\text{i.e.}\quad a_{r}(k)=\frac{1}{2\pi}\int_{-\pi}^{\pi}\sigma(\theta)^{r}e^{{\rm i}k\theta}\,d\theta$
(the integral is understood in the distributional sense whenever $\sigma^{r}\notin L^{1}$). Then for $f\in\mathcal{S}^{\prime}({\mathbb{Z}})$,
$\widehat{a_{r}*f}=\widehat{a_{r}}\,\widehat{f}=\vartheta^{r}\,\widehat{u}=\widehat{\Delta_{{\mathbb{Z}}}^{\,r}f}$.
Injectivity of $\mathcal{F}$ on $\mathcal{S}^{\prime}({\mathbb{Z}})$ yields $\Delta_{{\mathbb{Z}}}^{\,r}f=a_{r}*u$, i.e.
$(\Delta_{{\mathbb{Z}}}^{\,r}f)(n)=\sum_{k\in{\mathbb{Z}}}a_{r}(k)\,f(n-k),$
again in the distributional sense. This proves the proposition. ∎
@@
  • Integer $r\in\mathbb{N}^{*}$: $a_{r}$ has finite support (a local stencil), so $\Delta_{\mathbb{Z}}^{\,r}$ is a finite linear combination of shifts.
@@
    $a_{r}(k)=c_{r}\,|k|^{-1-2r}+\mathcal{O}\bigl{(}|k|^{-3-2r}\bigr{)},$
    for some $c_{r}\neq 0$.
@@
  • Negative $r<0$: writing r=sr=-s with $s>0$, the integrand behaves like $|\theta|^{2r}$ near $\theta=0$. For $\frac{1}{2}<r<0$ (i.e. $0<s\frac{1}{2}$) the integral defining $a_{r}(k)$ converges absolutely and $a_{r}$ is a function with
    $a_{-s}(k)=c_{-s}\,|k|^{\,2s-1}+\mathcal{O}\bigl{(}|k|^{\,2s-3}\bigr{)}\qquad(|k|\to\infty).$
    For $r\leq-\frac{1}{2}$ ($s\geq\tfrac{1}{2}$), $a_{r}$ must be understood as a tempered distribution (the symbol is non-integrable at $\theta=0$). In all cases $r<0$, one has $a_{r}\notin\ell^{1}(\mathbb{Z})$.
@@ 2.2. Anisotropic Fractional Powers of Translation-Invariant Laplacians on $\mathbb{Z}^{d}$
$\mathcal{D}\!\left(\Delta_{\mathbb{Z}^{d}}^{\vec{\mathbf{r}}}\right)=\bigcap_{j=1}^{d}\Bigl{(}\mathrm{id}^{\otimes(j-1)}\otimes\mathcal{D}(\Delta_{\mathbb{Z}}^{r_{j}})\otimes\mathrm{id}^{\otimes(d-j)}\Bigr{)}.$
@@
up to identification via the discrete Fourier transform.
@@ 2.3.1. Localized regularity and Mourre estimate
where $\mathrm{sign}(r_{j})=1$ if $r_{j}\geq 0,\ -1$ if $r_{j}<0$. Each $\mathrm{id}_{\ell^{2}({\mathbb{Z}}^{d})}$ is essentially self-adjoint on $\mathcal{C}_{c}(\mathbb{Z}^{d})$. We refer to [GeGo] and [Mic, Lemma 5.7] for the essential self-adjointness and [GeGo, Lemma 3.1] for the domain.
@@
Since $\Delta_{{\mathbb{Z}},j}$ is bounded self-adjoint with $\sigma(\Delta_{j})\subset[0,4]$, the fractional power $\Delta_{{\mathbb{Z}},j}^{r_{j}}$ is defined via the functional calculus, and $\lambda\mapsto\lambda^{r_{j}}$ is analytic on $(0,4)$. For polynomials, the commutator identity $\Delta_{j}^{\ell},\mathrm{i}A_{\mathbb{Z},r_{j}}]=\ell\,\Delta_{{\mathbb{Z}},j}^{\ell-1}[\Delta_{j},\mathrm{i}A_{\mathbb{Z},r_{j}}]$ holds, and extends to fractional powers via functional calculus (see [AmBoGe, Prop. 5.1.5]). Substituting $\Delta_{{\mathbb{Z}},j},\mathrm{i}A_{\mathbb{Z},r_{j}}]=\mathrm{sign}(r_{j})(4-\Delta_{{\mathbb{Z}},j})$ gives
$\Delta_{{\mathbb{Z}},j}^{r_{j}},\mathrm{i}A_{\mathbb{Z},r_{j}}]=|r_{j}|\,(4-\Delta_{{\mathbb{Z}},j})\,\Delta_{{\mathbb{Z}},j}^{r_{j}-1}.$
Summing over $j$ yields the quadratic form expression of $\Delta_{{\mathbb{Z}}^{d}}^{\vec{\mathbf{r}}},\mathrm{i}A_{{\mathbb{Z}}^{d},\vec{\mathbf{r}}}]$.
The symbol of $\Delta_{{\mathbb{Z}}^{d}}^{\vec{\mathbf{r}}},\mathrm{i}A_{{\mathbb{Z}}^{d},\vec{\mathbf{r}}}]$ in Fourier variables is strictly positive on $\mathcal{I}$ by construction, so the standard Mourre estimate follows from [AmBoGe, Thm. 4.4.2] without compact remainder terms. ∎
@@
For $j\in\{1,\dots,d\}$ and $r_{j}\in{\mathbb{R}}\setminus\{0\}$, define

### v2

@@ 1. Introduction and main results
Classical discrete Laplacians capture only short-range, nearest-neighbour interactions. Modern models of anomalous diffusion and long-range transport, however, require operators whose influence decays algebraically rather than being compactly supported. This motivates the study of fractional powers of the discrete Laplacians [JoKoPaSe, BaDeJiLi, Kw, Or, Or1] and their magnetic perturbations [FaZh]: they naturally describe long-range effects on graphs [TaZa], generate nonlocal transport phenomena [ScSo], and provide a bridge between discrete models and continuous nonlocal PDEs.
@@
The second approach is a Series or kernel representations. Explicit expansions in terms of lattice shifts or finite differences reveal how distant sites influence each other [Kw], a key point for both numerics and decay analysis. In the translation-invariant case of the infinite lattice $\mathbb{Z}^{d}$, the situation is considerably simpler. The discrete Laplacian is diagonalized by the Fourier transform, with symbol
@@
A parallel line of research initiated by Ortigueira has focused on developing a discrete-time fractional calculus. In [Or], the distinction between Riesz potentials and fractional Laplacians was emphasized in the signal-processing context, while his survey [Or1] traced the origins and evolutions of discrete fractional difference operators and introduced new formalisms. More recently, in [Or2] he proposed a multidimensional DTFT-based construction of fractional derivatives, essentially in the isotropic case $\alpha>0$. Our contribution is of a complementary nature: instead of a transform-based calculus, we construct anisotropic fractional Laplacians $\Delta_{{\mathbb{Z}}^{d}}^{\vec{r}}$ with general exponents $r_{j}\in{\mathbb{R}}$ as self-adjoint operators on $\ell^{2}({\mathbb{Z}}^{d})$, analyze their spectra and domains (including negative orders), and develop a full operator-theoretic framework leading to Mourre estimates, limiting absorption principles, and scattering. This technical shift—from signal-processing motivations to spectral and dynamical analysis—is at the core of the novelty of the present work, placing it at the crossroads between harmonic analysis, spectral theory, and mathematical physics. .
@@
In what follows we focus on the fractional discrete anisotropic Laplacian $\Delta^{\vec{\mathbf{r}}}_{{\mathbb{Z}}^{d}}$ acting on $\ell^{2}({\mathbb{Z}}^{d})$, where $\vec{\mathrm{r}}=(r_{1},\dots,r_{d})\neq 0$:
$\Delta^{\vec{\mathrm{r}}}_{{\mathbb{Z}}^{d}}:=\sum_{j=1}^{d}\mathrm{id}^{\otimes(j-1)}\otimes\Delta_{{\mathbb{Z}}}^{\,r_{j}}\otimes\mathrm{id}^{\otimes(d-j)}.$
@@
$H:=\Delta^{\vec{\mathrm{r}}}_{{\mathbb{Z}}^{d}}+W(Q),$
@@
$\sigma\!\left(\Delta_{{\mathbb{Z}}^{d}}^{\vec{\mathrm{r}}}\right)=\Bigl{\{}\ \sum_{j=1}^{d}\lambda_{j}\ \Big{|}\ \lambda_{j}\in\sigma\!\left(\Delta_{{\mathbb{Z}}}^{\,r_{j}}\right)\ \Bigr{\}},$
@@
$\operatorname{Thr}_{\mathrm{fin}}\!\bigl{(}H_{{\mathbb{Z}}^{d}}^{\vec{\mathrm{r}}}\bigr{)}=\Bigl{\{}\ \sum_{j\in N}4^{\,r_{j}}+\sum_{j\in P}\epsilon_{j}\,4^{\,r_{j}}\ :\ \epsilon_{j}\in\{0,1\}\ \Bigr{\}},$
@@
Let $H=\Delta_{{\mathbb{Z}}^{d}}^{\vec{\mathrm{r}}}+W(Q)$ with real $W$ satisfying (H0) and (H1). Then $H\in\mathcal{C}^{1,1}_{\mathrm{loc}}(A_{{\mathbb{Z}}^{d}})$ on each $\mathcal{I}\Subset\sigma(H_{0})^{\circ}\setminus\operatorname{Thr}(\Delta_{{\mathbb{Z}}^{d}}^{\vec{\mathrm{r}}})$, and localized Mourre theory yields:
@@
This work develops a localized (nonlocal) Mourre theory for fractional discrete Laplacians $\Delta^{\vec{\mathrm{r}}}_{{\mathbb{Z}}^{d}}$ with anisotropic orders $\vec{\mathrm{r}}\neq 0$. Within this unified framework we prove, on compact energy windows away from thresholds, the Limiting Absorption Principle and asymptotic completeness for fractional difference operators. To the best of our knowledge, this is the first rigorous Mourre/LAP/scattering treatment for fractional discrete Laplacians.
@@ 2.1. Fractional powers of discrete Laplacians in the unidimensional setting
The (forward) shift operator $U$ on $\ell^{2}(\mathbb{Z})$ is defined by
$(Uf)(n):=f(n-1),\qquad(U^{*}f)(n):=f(n+1).$
Then $U$ is unitary and $UU^{*}=U^{*}U=\mathrm{id}_{\ell^{2}(\mathbb{Z})}$.
We also denote by $Q$ the position operator on $\ell^{2}(\mathbb{Z})$ defined by $(Qf)(n)=nf(n)$. For any $s\in\mathbb{R}$, we define the weighted space $\ell^{2}_{s}(\mathbb{Z}):=\{f\in\ell^{2}(\mathbb{Z})\mid\|\langle Q\rangle^{s}f\|<\infty\}$, where $\langle Q\rangle:=(1+Q^{2})^{1/2}$. Let $\Delta_{\mathbb{Z}}:=2\mathrm{id}-U-U^{*}$ denote the standard discrete Laplacian on $\ell^{2}(\mathbb{Z})$. This operator is self-adjoint and nonnegative on $\ell^{2}(\mathbb{Z})$. To extend the notion of fractional powers $\Delta^{r}_{\mathbb{Z}}$ for any real $r$, we use the discrete Fourier transform framework and the Borel functional calculus; this defines $\Delta^{r}_{\mathbb{Z}}$ also for $r<0$ (as an unbounded operator with appropriate domain).
@@
$\widehat{\Delta_{{\mathbb{Z}}}^{\,r}u}(\theta)=\bigl{(}2(1-\cos\theta)\bigr{)}^{r}\,\widehat{u}(\theta),\qquad(\Delta_{{\mathbb{Z}}}^{\,r}u)(n)=\sum_{k\in{\mathbb{Z}}}a_{r}(k)\,u(n-k),$
@@
$a_{r}(k)=\frac{1}{2\pi}\int_{-\pi}^{\pi}\bigl{(}2(1-\cos\theta)\bigr{)}^{r}e^{\mathrm{i}k\theta}\,d\theta.$
@@
$(\mathcal{F}u)(\theta)=\frac{1}{\sqrt{2\pi}}\sum_{n\in{\mathbb{Z}}}u(n)e^{-\mathrm{i}n\theta}.$
@@
$(\mathcal{F}U\mathcal{F}^{*}f)(\theta)=e^{-\mathrm{i}\theta}f(\theta),\qquad(\mathcal{F}U^{*}\mathcal{F}^{*}f)(\theta)=e^{\mathrm{i}\theta}f(\theta),$
@@
By the Borel functional calculus, $\mathcal{F}\,\Delta_{{\mathbb{Z}}}^{\,r}\,\mathcal{F}^{*}=M_{\vartheta^{r}}$, whence the multiplier identity on $\ell^{2}({\mathbb{Z}})$; the extension to $\mathcal{S}^{\prime}({\mathbb{Z}})$ follows by density/duality. Define $a_{r}$ by $\widehat{a_{r}}=\vartheta^{r}$ (as a tempered distribution when needed). Then $\widehat{a_{r}*u}=\widehat{a_{r}}\,\widehat{u}=\vartheta^{r}\widehat{u}=\widehat{\Delta_{\mathbb{Z}}^{\,r}u}$, and injectivity of $\mathcal{F}$ on $\mathcal{S}^{\prime}({\mathbb{Z}})$ yields $\Delta_{\mathbb{Z}}^{\,r}u=a_{r}*u$. Evenness and reality are clear from the integrand. ∎
@@
  • Integer $r\in\mathbb{N}^{*}$: $a_{r}$ has finite support (local stencil).
@@
    $a_{r}(k)=c_{r}\,|k|^{-1-2r}+\mathcal{O}\bigl{(}|k|^{-3-2r}\bigr{)},\qquad c_{r}\neq 0.$
@@
  • Negative $r<0$: write r=sr=-s with $s>0$. Near $\theta=0$ the integrand behaves like $|\theta|^{2r}$. For $\tfrac{1}{2}<r<0$ (i.e. $0<s<\tfrac{1}{2}$) the integral defining $a_{r}(k)$ converge absolutely and
    $a_{-s}(k)=c_{-s}\,|k|^{\,2s-1}+\mathcal{O}\bigl{(}|k|^{\,2s-3}\bigr{)}\quad(|k|\to\infty).$
    For $r\leq-\tfrac{1}{2}$ ($s\geq\tfrac{1}{2}$), $a_{r}$ is only a tempered distribution. In all cases $r<0$, $a_{r}\notin\ell^{1}({\mathbb{Z}})$.
@@ 2.2. Anisotropic Fractional Powers of Translation-Invariant Laplacians on $\mathbb{Z}^{d}$
$\mathcal{D}\!\left(\Delta_{\mathbb{Z}^{d}}^{\vec{\mathrm{r}}}\right)=\bigcap_{j=1}^{d}\Bigl{(}\mathrm{id}^{\otimes(j-1)}\otimes\mathcal{D}(\Delta_{\mathbb{Z}}^{r_{j}})\otimes\mathrm{id}^{\otimes(d-j)}\Bigr{)}.$
@@
up to identification via the discrete Fourier transform. Anisotropic Sobolev scale (case $r_{j}>0$). Via Fourier, the form domain above is equivalent to the weighted space
$H^{\vec{\mathrm{r}}}({\mathbb{T}}^{d}):=\Big{\{}f\in L^{2}({\mathbb{T}}^{d}):\int_{{\mathbb{T}}^{d}}\Big{(}1+\sum_{j=1}^{d}\lambda_{j}(\theta)^{r_{j}}\Big{)}|\hat{f}(\theta)|^{2}\,d\theta<\infty\Big{\}}.$
For directions with $r_{j}<0$ we keep the domain description through $\mathcal{D}(\Delta_{{\mathbb{Z}}}^{r_{j}})$ and the intersection rule.
@@ 2.3.1. Localized regularity and Mourre estimate
where $\mathrm{sign}(r_{j})=1$ if $r_{j}\geq 0,\ -1$ if $r_{j}<0$. Each $A_{\mathbb{Z},r_{j}}$ is essentially self-adjoint on $\mathcal{C}_{c}(\mathbb{Z}^{d})$. We refer to [GeGo] and [Mic, Lemma 5.7] for the essential self-adjointness and [GeGo, Lemma 3.1] for the domain.
@@
Since $\Delta_{{\mathbb{Z}},j}$ is bounded self–adjoint with $\sigma(\Delta_{{\mathbb{Z}},j})=[0,4]$, the fractional power $\Delta_{{\mathbb{Z}},j}^{r_{j}}$ is defined by the functional calculus. For the function $f(\lambda)=\lambda^{r_{j}}$ one has $f^{\prime}(\lambda)=r_{j}\,\lambda^{r_{j}-1}$, which is bounded on every compact subinterval of $(0,4)$. Therefore, for any spectral localization $E_{\mathcal{I}}(\Delta_{{\mathbb{Z}},j})$ with $\mathcal{I}\Subset(0,4)$, the chain rule for commutators (see [AmBoGe, Prop. 5.1.5]) applies and yields
$\Delta_{{\mathbb{Z}},j}^{r_{j}},\mathrm{i}A_{\mathbb{Z},r_{j}}]=r_{j}\,\Delta_{{\mathbb{Z}},j}^{\,r_{j}-1}\,[\Delta_{{\mathbb{Z}},j},\mathrm{i}A_{\mathbb{Z},r_{j}}]\quad\text{on }\mathrm{Ran}E_{\mathcal{I}}(\Delta_{{\mathbb{Z}},j}).$
Since $\Delta_{{\mathbb{Z}},j},\mathrm{i}A_{\mathbb{Z},r_{j}}]=\mathrm{sign}(r_{j})(4-\Delta_{{\mathbb{Z}},j})$, it follows that
$\Delta_{{\mathbb{Z}},j}^{r_{j}},\mathrm{i}A_{\mathbb{Z},r_{j}}]=|r_{j}|\,(4-\Delta_{{\mathbb{Z}},j})\,\Delta_{{\mathbb{Z}},j}^{\,r_{j}-1}\quad\text{on }\mathrm{Ran}E_{\mathcal{I}}(\Delta_{{\mathbb{Z}},j}).$
@@
For $j\in\{1,\dots,d\}$ and $r_{j}\in{\mathbb{R}}\setminus\{0\}$, we define